why does (3-k)/4=(k+3)/2 equal -1? in the end it gives 0=0 so why is it -1?
Found 2 solutions by jorel1380, richard1234:
Answer by jorel1380(3719)

3-k/4=k+3/2
2(3-k)=4(k+3)
6-2k=4k+12
6k=-6
k=-1..

Answer by richard1234(7193)

If the equation was equivalent to 0 = 0, this would imply any value for k would work, which does not seem to be the case. Also, be sure to rephrase your question "why does ... equal -1" when you really mean "why does ... have a solution k = -1."

If we cross multiply, we get

2(3-k) = 4(k+3)

-2k + 6 = 4k + 12

-6k = 6

k = -1